# The Complexity of HyperQPTL

### v2

@@ Abstract
HyperQPTL and HyperQPTL+ are expressive specification languages for hyperproperties, i.e., properties that relate multiple executions of a system. Tight complexity bounds are known for HyperQPTL finite-state satisfiability and model-checking.
@@ 1 Introduction
Hyperproperties [3] are properties relating multiple executions of a system and have found applications in security and privacy, epistemic reasoning, and verification. Temporal logics have been introduced to express hyperproperties, e.g., $\mathrm{Hyper\-LTL}$ and $\mathrm{HyperCTL^{*}}$ [2] (which extend $\mathrm{LTL}$ and $\mathrm{CTL^{*}}$ with trace quantification), $\mathrm{Hyper^{2}LTL}$ [1] (second-order $\mathrm{Hyper\-LTL}$, which extends $\mathrm{Hyper\-LTL}$ with quantification over sets of traces), and many more. Here, we are interested in the most important verification problems:
  • Satisfiability: Given a sentence $\varphi$, does it have a model, i.e., is there a nonempty set $T$ of traces such that $T\models\varphi$?
@@
This work is part of a research program [6, 11, 4, 7, 8, 12] settling the complexity of these problems for hyperlogics. Most importantly for applications in verification, the model-checking problems for $\mathrm{Hyper\-LTL}$ and $\mathrm{HyperCTL^{*}}$ are decidable, albeit Tower-complete [6, 11, 10]. However, satisfiability is typically much harder. In fact, the satisfiability problems are typically highly undecidable, i.e., we measure their complexity by placing them in the arithmetic or analytic hierarchy, or even beyond: Intuitively, first-order arithmetic is predicate logic over the signature $(+,\cdot,<)$ where quantification ranges over natural numbers. Similarly, second-order arithmetic adds quantification over sets of natural numbers to first-order arithmetic while third-order arithmetic adds quantification over sets of sets of natural numbers to second-order quantification. Figure 1 gives an overview of the arithmetic, analytic, and “third” hierarchy, each spanned by the classes of languages definable by restricting the number of alternations of the highest-order quantifiers, i.e., $\Sigma_{n}^{0}$ contains languages definable by formulas of first-order arithmetic with n1n-1 quantifier alternations, starting with an existential one.
@@
Here, we study the logics $\mathrm{Hyper\-QPTL}$ and $\mathrm{Hyper\-QPTL^{+}}$ which extend $\mathrm{Hyper\-LTL}$ by quantification over propositions [11, 5], just as $\mathrm{QPTL}$ extends $\mathrm{LTL}$ with quantification over propositions [14]. $\mathrm{QPTL}$ is, unlike $\mathrm{LTL}$, able to express all $\omega$-regular properties while $\mathrm{Hyper\-QPTL}$ allows to express, e.g., promptness and epistemic properties, which are not expressible in $\mathrm{Hyper\-LTL}$ [5].
@@
Table 1 presents our results (in bold) as well as results for the related logics mentioned above. There, $\mathrm{Hyper^{2}LTL_{\mathrm{mm}}}$ and $\mathrm{lfp\text{-}{Hyper^{2}LTL_{\mathrm{mm}}}}$ are the fragments of second-order $\mathrm{Hyper\-LTL}$ obtained by restricting set quantification to minimal/maximal sets satisfying a guard formula respectively by restricting set quantification to least fixed points of $\mathrm{LTL}$-definable operators [1, 8]. [^1] Note that [1] and [8] use different names for these fragments. We follow [8]. Furthermore, $\mathrm{Hyper^{2}LTL}$ comes with two semantics, standard and closed-world (CW). In most cases, the choice of semantics does not influence the complexity of the verification problems. The only known exception is the satisfiability problem, which is $\Sigma_{1}^{2}$-complete for standard semantics, but only $\Sigma_{1}^{1}$-complete for closed-world semantics [8, 12].
Table 1: List of our results (in bold) and comparison to related logics. “T2A-equivalent” stands for “equivalent to truth in second-order arithmetic”, “T3A-equivalent” for “equivalent to truth in third-order arithmetic”.
Logic Satisfiability Finite-state satisfiability Model-checking
$\mathrm{LTL}$ PSpace-complete [15] PSpace-complete [15] PSpace-complete [15]
$\mathrm{Hyper\-LTL}$ $\Sigma_{1}^{1}$-complete [7] $\Sigma_{1}^{0}$-complete [4] Tower-complete [6, 11, 10]
$\mathrm{Hyper^{2}LTL}$ T3A-equivalent [8] T3A-equivalent [8] T3A-equivalent [8]
$\mathrm{Hyper^{2}LTL_{\mathrm{mm}}}$ T3A-equivalent [8] T3A-equivalent [8] T3A-equivalent [8]
$\mathrm{lfp\text{-}{Hyper^{2}LTL_{\mathrm{mm}}}}$ $\Sigma_{1}^{2}$-complete [12]/ T2A-equivalent [12] T2A-equivalent [12]
$\Sigma_{1}^{1}$-complete (CW) [8]
$\mathrm{Hyper\-QPTL}$ $\Sigma^{2}_{1}$-complete $\Sigma_{1}^{0}$-complete [11] Tower-complete [11]
$\mathrm{Hyper\-QPTL^{+}}$ T3A-equivalent T3A-equivalent T3A-equivalent
@@ 2 Preliminaries
We denote the nonnegative integers by $\mathbb{N}$.
2.1 Words, Traces, and Transition Systems
An alphabet is a nonempty finite set. The set of infinite words over an alphabet $\Sigma$ is denoted by $\Sigma^{\omega}$. Let $\mathrm{AP}$ be a nonempty finite set of atomic propositions. A trace over $\mathrm{AP}$ is an infinite word over the alphabet $2^{\mathrm{AP}}$. Given a subset $\mathrm{AP}^{\prime}\subseteq\mathrm{AP}$, the $\mathrm{AP}^{\prime}$-projection of a trace $t(0)t(1)t(2)\cdots$ over $\mathrm{AP}$ is the trace $(t(0)\cap\mathrm{AP}^{\prime})(t(1)\cap\mathrm{AP}^{\prime})(t(2)\cap\mathrm{AP}^{\prime})\cdots\in(2^{\mathrm{AP}^{\prime}})^{\omega}$. The $\mathrm{AP}^{\prime}$-projection of $T\subseteq(2^{\mathrm{AP}})^{\omega}$ is defined as the set of $\mathrm{AP}$-projections of traces in $T$. We write $t=_{\mathrm{AP}^{\prime}}t^{\prime}$ ($T=_{\mathrm{AP}^{\prime}}T^{\prime}$) if the $\mathrm{AP}^{\prime}$-projections of $t$ and $t^{\prime}$ ($T$ and $T^{\prime}$) are equal. Now, let $\mathrm{AP}$ and $\mathrm{AP}^{\prime}$ be two disjoint sets, let $t$ be a trace over $\mathrm{AP}$, and let $t^{\prime}$ be a trace over $\mathrm{AP}^{\prime}$. Then, we define $t^{\smallfrown}t^{\prime}$ as the pointwise union of $t$ and $t^{\prime}$, i.e., $t^{\smallfrown}t^{\prime}$ is the trace over $\mathrm{AP}\cup\mathrm{AP}^{\prime}$ defined as $(t(0)\cup t^{\prime}(0))(t(1)\cup t^{\prime}(1))(t(2)\cup t^{\prime}(2))\cdots$.
@@ 2.2 Arithmetic
Truth in third-order arithmetic is the following problem: given a sentence $\varphi$ of third-order arithmetic, does (,+,,<,)($\mathbb{N}$$\cdot,$$\in$) satisfy $\varphi$ (written $(\mathbb{N},+,\cdot,<,\in)\models\varphi$ as usual)? Furthermore, $\Sigma^{2}_{1}$ contains the sets of the form
$\{n\in\mathbb{N}\mid(\mathbb{N},+,\cdot,<,\in)\models\varphi(n)\},$
where $\varphi$ is a formula of the form $\exists\mathcal{Y}_{1}.\ \ldots\exists\mathcal{Y}_{k}.\ \psi(x,\mathcal{Y}_{1},\ldots,\mathcal{Y}_{k},)$ where the $\mathcal{Y}_{j}$ are third-order variables and $\psi$ is a formula of arithmetic containing only first- and second-order quantifiers.
@@ 3.1 Syntax and Semantics
where p and q range over $\mathrm{AP}$ and where $\pi$ ranges over $\mathcal{V}$. Note that there are two types of atomic formulas, i.e., propositions labeled by traces on which they are evaluated ($\texttt{p}_{\pi}$ with $\texttt{p}\in\mathrm{AP}$ and $\pi\in\mathcal{V}$) and unlabeled propositions ($\texttt{q}\in\mathrm{AP}$). [^2] We use different letters for the propositions in these cases, but let us stress again that both p and q are propositions in $\mathrm{AP}$. A formula is a sentence, if every occurrence of an atomic formula $\texttt{p}_{\pi}$ is in the scope of a quantifier binding $\pi$ and every occurrence of an atomic formula q is in the scope of a quantifier binding q. Finally, note that the only temporal operators we have in the syntax are next ($\mathop{\mathbf{X}\vphantom{a}}\nolimits$) and eventually ($\mathop{\mathbf{F}\vphantom{a}}\nolimits$), as the other temporal operators like always ($\mathop{\mathbf{G}\vphantom{a}}\nolimits$) and until $(\operatorname{\mathbf{U}})$ are syntactic sugar in $\mathrm{Hyper\-QPTL}$ [9]. Hence, we will use them freely in the following, just as we use conjunction, implication, and equivalence.
A trace assignment is a partial mapping $\Pi\colon\mathcal{V}\rightarrow(2^{\mathrm{AP}})^{\omega}$. Given a variable $\pi\in\mathcal{V}$, and a trace $t$, we denote by $\Pi[\pi\mapsto t]$ the trace assignment that coincides with $\Pi$ on all variables but $\pi$, which is mapped to $t$. It remains to capture the semantics of quantification of propositions $\texttt{q}\in\mathrm{AP}$. Let $t\in(2^{\mathrm{AP}})^{\omega}$ be a trace over $\mathrm{AP}$ and $t_{\texttt{q}}\in(2^{\{\texttt{q}\}})^{\omega}$ be a trace over $\{\texttt{q}\}$. We define the trace $t[\texttt{q}\mapsto t_{\texttt{q}}]=t^{\prime}{}^{\smallfrown}t_{\texttt{q}}$, where $t^{\prime}$ is the $(\mathrm{AP}\setminus\{\texttt{q}\})$-projection of $t$: Intuitively, the occurrences of q in $t$ are replaced according to $t_{\texttt{q}}$. We lift this to sets $T$ of traces by defining $T[\texttt{q}\mapsto t_{\texttt{q}}]=\{t[\texttt{q}\mapsto t_{\texttt{q}}]\mid t\in T\}$. Note that all traces in $T[\texttt{q}\mapsto t_{\texttt{q}}]$ have the same $\{\texttt{q}\}$-projection, which is $t_{\texttt{q}}$.
@@
  • $T,\Pi,i\models\mathop{\mathbf{F}\vphantom{a}}\nolimits\psi$ if there is an $i^{\prime}\geq i$ such that $T,\Pi,i^{\prime}\models\psi$,
  • $T,\Pi,i\models\exists\pi.\,\varphi$ if there exists a trace $t\in T$ such that $T,\Pi[\pi\mapsto t],i\models\varphi$,
  • $T,\Pi,i\models\forall\pi.\ \varphi$ if for all traces $t\in T$ we have $T,\Pi[\pi\mapsto t],i\models\varphi$,
  • $T,\Pi,i\models\exists\texttt{q}.\ \varphi$ if there exists a trace $t_{\texttt{q}}\in(2^{\{\texttt{q}\}})^{\omega}$ such that $T[\texttt{q}\mapsto t_{\texttt{q}}],\Pi,i\models\varphi$, and
  • $T,\Pi,i\models\forall\texttt{q}.\ \varphi$ if for all traces $t_{\texttt{q}}\in(2^{\{\texttt{q}\}})^{\omega}$ we have $T[\texttt{q}\mapsto t_{\texttt{q}}],\Pi,i\models\varphi$.
@@
While it is known that $\mathrm{Hyper\-QPTL}$ finite-state satisfiability is $\Sigma^{0}_{1}$-complete [11], i.e., complete for the recursively-enumerable languages, and $\mathrm{Hyper\-QPTL}$ model-checking is Tower-complete [11], the exact complexity of $\mathrm{Hyper\-QPTL}$ satisfiability is open: it is only known to be $\Sigma_{1}^{1}$-hard, as satisfiability is already $\Sigma_{1}^{1}$-complete for its fragment $\mathrm{Hyper\-LTL}$ [7].
@@ 3.2 HyperQPTL Satisfiability
In this subsection, we study the $\mathrm{Hyper\-QPTL}$ satisfiability problem, which asks whether, for a given $\mathrm{Hyper\-QPTL}$ sentence $\varphi$, there is a nonempty $T$ such that $T\models\varphi$. We show that it is $\Sigma^{2}_{1}$-complete: intuitively, quantification over propositions q (which, on the semantical level, corresponds to quantification of traces over $2^{\{\texttt{q}\}}$) is equivalent to quantification over subsets of $\mathbb{N}$: a trace $t\in(2^{\{\texttt{q}\}})^{\omega}$ is encoded by the set $\{n\in\mathbb{N}\mid\texttt{q}\in t(n)\}$ and vice versa. Furthermore, existential quantification over sets of sets of natural numbers is equivalent to the selection of a set of traces at the meta-level of the satisfiability problem.
For the lower bound, we need to be able to enforce that the model of a $\mathrm{Hyper\-QPTL}$ sentence contains enough traces to encode all subsets of $\mathbb{N}$. As usual, we use $\mathfrak{c}$ to denote the cardinality of the continuum, or equivalently, the cardinality of $(2^{\mathrm{AP}})^{\omega}$ for each finite $\mathrm{AP}$ and the cardinality of $2^{\mathbb{N}}$. As models of $\mathrm{Hyper\-QPTL}$ sentences are sets of traces, $\mathfrak{c}$ is a trivial upper bound on the size of models. It is straightforward to show that this upper bound is tight, which also implies that there are sentences whose models allow us to encode all subsets of $\mathbb{N}$.
@@
We need to show that every language in $\Sigma^{2}_{1}$ is polynomial-time reducible to $\mathrm{Hyper\-QPTL}$ satisfiability. To this end, we first present a polynomial-time computable function $f$ mapping sentences of the form $\varphi=\exists\mathcal{Y}_{1}.\ \ldots\exists\mathcal{Y}_{k}.\ \psi$, with third-order variables $\mathcal{Y}_{j}$ and with $\psi$ only containing first-order and second-order quantifiers, to $\mathrm{Hyper\-QPTL}$ sentences $f(\varphi)$ such that $(\mathbb{N},+,\cdot,<,\in)\models\varphi$ if and only if $f(\varphi)$ is satisfiable. To this end, we fix $\mathrm{AP}=\{\texttt{x},\texttt{q},\texttt{m}_{1},\ldots,\texttt{m}_{k},\texttt{arg1},\texttt{arg2},\texttt{res},\texttt{add},\texttt{mult}\}$. We explain the roles of the different propositions along the way.
@@
$\theta_{\mathit{cons}}=\bigwedge_{j=1}^{k}\forall\pi.\ \mathop{\mathbf{X}\vphantom{a}}\nolimits\mathop{\mathbf{G}\vphantom{a}}\nolimits\neg(\texttt{m}_{j})_{\pi}\wedge\forall\pi^{\prime}.\ \Big(\big(\mathop{\mathbf{G}\vphantom{a}}\nolimits(\texttt{x}_{\pi}\leftrightarrow\texttt{x}_{\pi^{\prime}})\big)\rightarrow\big((\texttt{m}_{j})_{\pi}\leftrightarrow(\texttt{m}_{j})_{\pi^{\prime}}\big)\Big)$
to express consistency: the first part expresses that only the first position of a trace may (but must not) be labeled with markers and the second part expresses that when two traces coincide on their truth values of x at every position (i.e., they encode the same set), then they have the same marking.
Thus, every model $T$ of $\theta_{\mathit{all}}\wedge\theta_{\mathit{cons}}$ encodes every subset of $\mathbb{N}$ as a trace in $T$ and also $k$ sets of subsets of $\mathbb{N}$ via the (consistent) marking of traces. Thus, the existential quantification over a model in the satisfiability problem can simulate the existential third-order quantification of the $\mathcal{Y}_{j}$.
Finally, Fortin et al. showed that addition and multiplication can be “implemented” in $\mathrm{Hyper\-LTL}$ [7]: Let $T_{(+,\cdot)}$ be the set of all traces $t\in(2^{\mathrm{AP}})^{\omega}$ such that
  • there are unique $n_{1},n_{2},n_{3}\in\mathbb{N}$ with $\texttt{arg1}\in t(n_{1})$, $\texttt{arg2}\in t(n_{2})$, and $\texttt{res}\in t(n_{3})$, and
  • either
@@
Now, we define for $\varphi=\exists\mathcal{Y}_{1}.\ \ldots\exists\mathcal{Y}_{k}.\ \psi$
@@
  • For second-order variables $Y$, ${\mathit{hyp}}(\exists Y.\ \psi)=\exists\pi_{Y}.\ {\mathit{hyp}}(\psi)$.
  • For second-order variables $Y$, ${\mathit{hyp}}(\forall Y.\ \psi)=\forall\pi_{Y}.\ {\mathit{hyp}}(\psi)$.
  • For first-order variables $y$,
    ${\mathit{hyp}}(\exists y.\ \psi)=\exists\pi_{y}.\ ((\neg\texttt{x}_{\pi_{y}})\operatorname{\mathbf{U}}(\texttt{x}_{\pi_{y}}\wedge\mathop{\mathbf{X}\vphantom{a}}\nolimits\mathop{\mathbf{G}\vphantom{a}}\nolimits\neg\texttt{x}_{\pi_{y}}))\land{\mathit{hyp}}(\psi)$,
  • For first-order variables $y$,
    ${\mathit{hyp}}(\forall y.\ \psi)=\forall\pi_{y}.\ ((\neg\texttt{x}_{\pi_{y}})\operatorname{\mathbf{U}}(\texttt{x}_{\pi_{y}}\wedge\mathop{\mathbf{X}\vphantom{a}}\nolimits\mathop{\mathbf{G}\vphantom{a}}\nolimits\neg\texttt{x}_{\pi_{y}}))\rightarrow{\mathit{hyp}}(\psi)$,

### v3

@@ Abstract
HyperQPTL and HyperQPTL+ are expressive specification languages for hyperproperties, properties that relate multiple executions of a system. Tight complexity bounds are known for HyperQPTL finite-state satisfiability and model-checking.
@@ 1 Introduction
Hyperproperties [3] are properties relating multiple executions of a system and have found applications in security and privacy, epistemic reasoning, and verification. Temporal logics have been introduced to express hyperproperties, e.g., $\mathrm{Hyper\-LTL}$ and $\mathrm{HyperCTL^{*}}$ [2] (which extend $\mathrm{LTL}$ and $\mathrm{CTL^{*}}$ with trace quantification), $\mathrm{Hyper^{2}LTL}$ [1] (second-order $\mathrm{Hyper\-LTL}$, which extends $\mathrm{Hyper\-LTL}$ with quantification over sets of traces), and many more. Here, we focus on the most important verification problems:
  • Satisfiability: Given a sentence $\varphi$, is there a nonempty set $T$ of traces such that $T\models\varphi$?
@@
This work is part of a research program [6, 11, 5, 7, 8, 12, 13] settling the complexity of these problems for hyperlogics. Most importantly for applications in verification, the model-checking problems for $\mathrm{Hyper\-LTL}$ and $\mathrm{HyperCTL^{*}}$ are decidable, albeit Tower-complete [6, 11, 10]. However, satisfiability is typically much harder. In fact, the satisfiability problems are typically highly undecidable, i.e., we measure their complexity by placing them in the arithmetic or analytic hierarchy, or even beyond: Intuitively, first-order arithmetic is predicate logic over the signature $(+,\cdot,<)$ where quantification ranges over natural numbers. Similarly, second-order arithmetic adds quantification over sets of natural numbers to first-order arithmetic while third-order arithmetic adds quantification over sets of sets of natural numbers to second-order quantification. Figure 1 depicts the arithmetic, analytic, and “third” hierarchy, each spanned by the classes of languages definable by restricting the number of alternations of the highest-order quantifiers, e.g., $\Sigma_{n}^{0}$ contains languages definable by formulas of first-order arithmetic with n1n-1 quantifier alternations, starting with an existential one.
@@
Here, we study the logics $\mathrm{Hyper\-QPTL}$ and $\mathrm{Hyper\-QPTL^{+}}$ which extend $\mathrm{Hyper\-LTL}$ by quantification over propositions [11, 4], just as $\mathrm{QPTL}$ extends $\mathrm{LTL}$ with quantification over propositions [15]. $\mathrm{QPTL}$ is, unlike $\mathrm{LTL}$, able to express all $\omega$-regular properties while $\mathrm{Hyper\-QPTL}$ is, unlike $\mathrm{Hyper\-LTL}$, able to express, e.g., promptness and epistemic properties [4].
@@ 2 Preliminaries
We denote the nonnegative integers by $\mathbb{N}$. An alphabet is a nonempty finite set. The set of infinite words over an alphabet $\Sigma$ is denoted by $\Sigma^{\omega}$. Let $\mathrm{AP}$ be a nonempty finite set of atomic propositions. A trace over $\mathrm{AP}$ is an infinite word over the alphabet $2^{\mathrm{AP}}$. Given a subset $\mathrm{AP}^{\prime}\subseteq\mathrm{AP}$, the $\mathrm{AP}^{\prime}$-projection of a trace $t(0)t(1)t(2)\cdots$ over $\mathrm{AP}$ is the trace $(t(0)\cap\mathrm{AP}^{\prime})(t(1)\cap\mathrm{AP}^{\prime})(t(2)\cap\mathrm{AP}^{\prime})\cdots\in(2^{\mathrm{AP}^{\prime}})^{\omega}$. The $\mathrm{AP}^{\prime}$-projection of $T\subseteq(2^{\mathrm{AP}})^{\omega}$ is defined as the set of $\mathrm{AP}^{\prime}$-projections of traces in $T$. We write $t_{0}=_{\mathrm{AP}^{\prime}}t_{1}$ ($T_{0}=_{\mathrm{AP}^{\prime}}T_{1}$) if the $\mathrm{AP}^{\prime}$-projections of $t_{0}$ and $t_{1}$ ($T_{0}$ and $T_{1}$) are equal. Now, let $\mathrm{AP}$ and $\mathrm{AP}^{\prime}$ be two disjoint sets, let $t$ be a trace over $\mathrm{AP}$, and let $t^{\prime}$ be a trace over $\mathrm{AP}^{\prime}$. Then, we define $t^{\smallfrown}t^{\prime}$ as the pointwise union of $t$ and $t^{\prime}$, i.e., $t^{\smallfrown}t^{\prime}$ is the trace over $\mathrm{AP}\cup\mathrm{AP}^{\prime}$ defined as $(t(0)\cup t^{\prime}(0))(t(1)\cup t^{\prime}(1))(t(2)\cup t^{\prime}(2))\cdots$.
@@
Truth in third-order arithmetic is the following problem: given a sentence $\varphi$ of third-order arithmetic, does (,+,,<,)($\mathbb{N}$$\cdot,$$\in$) satisfy $\varphi$ (written $(\mathbb{N},+,\cdot,<,\in)\models\varphi$ as usual)? Furthermore, $\Sigma^{2}_{1}$ contains the sets of the form $\{n\in\mathbb{N}\mid(\mathbb{N},+,\cdot,<,\in)\models\varphi(n)\},$ where $\varphi$ is a formula of the form $\exists\mathcal{Y}_{1}\ldots\exists\mathcal{Y}_{k}.\ \psi(x,\mathcal{Y}_{1},\ldots,\mathcal{Y}_{k},)$ where the $\mathcal{Y}_{j}$ are third-order variables and $\psi$ is a formula of arithmetic containing only first- and second-order quantifiers.
@@ 3.1 Syntax and Semantics
where p and q range over $\mathrm{AP}$ and where $\pi$ ranges over $\mathcal{V}$. Note that there are two types of atomic formulas, i.e., propositions labeled by traces on which they are evaluated ($\texttt{p}_{\pi}$ with $\texttt{p}\in\mathrm{AP}$ and $\pi\in\mathcal{V}$) and unlabeled propositions ($\texttt{q}\in\mathrm{AP}$). We use different letters for the propositions in these cases, but let us stress again that both p and q are propositions in $\mathrm{AP}$. A formula is a sentence, if every occurrence of an atomic formula $\texttt{p}_{\pi}$ is in the scope of a quantifier binding $\pi$ and every occurrence of an atomic formula q is in the scope of a quantifier binding q. Finally, note that the only temporal operators we have in the syntax are next ($\mathop{\mathbf{X}\vphantom{a}}\nolimits$) and eventually ($\mathop{\mathbf{F}\vphantom{a}}\nolimits$), as the other temporal operators like always ($\mathop{\mathbf{G}\vphantom{a}}\nolimits$) and until $(\operatorname{\mathbf{U}})$ are syntactic sugar in $\mathrm{Hyper\-QPTL}$ [9]. Hence, we use them freely in the following, just as conjunction, implication, and equivalence.
A trace assignment is a partial mapping $\Pi\colon\mathcal{V}\rightarrow(2^{\mathrm{AP}})^{\omega}$. Given a variable $\pi\in\mathcal{V}$, and a trace $t$, we denote by $\Pi[\pi\mapsto t]$ the trace assignment that coincides with $\Pi$ on all variables but $\pi$, which is mapped to $t$. Let $\texttt{q}\in\mathrm{AP}$, let $t\in(2^{\mathrm{AP}})^{\omega}$ be a trace over $\mathrm{AP}$, and let $t_{\texttt{q}}\in(2^{\{\texttt{q}\}})^{\omega}$ be a trace over $\{\texttt{q}\}$. We define the trace $t[\texttt{q}\mapsto t_{\texttt{q}}]=t^{\prime}{}^{\smallfrown}t_{\texttt{q}}$, where $t^{\prime}$ is the $(\mathrm{AP}\setminus\{\texttt{q}\})$-projection of $t$: Intuitively, the occurrences of q in $t$ are replaced according to $t_{\texttt{q}}$. We lift this to sets $T$ of traces by defining $T[\texttt{q}\mapsto t_{\texttt{q}}]=\{t[\texttt{q}\mapsto t_{\texttt{q}}]\mid t\in T\}$. All traces in $T[\texttt{q}\mapsto t_{\texttt{q}}]$ have the same $\{\texttt{q}\}$-projection, i.e., $t_{\texttt{q}}$.
@@
  • $T,\Pi,i\models\mathop{\mathbf{F}\vphantom{a}}\nolimits\psi$ if there is an $i^{\prime}\geq i$ with $T,\Pi,i^{\prime}\models\psi$,
  • $T,\Pi,i\models\exists\pi.\,\varphi$ if there exists a $t\in T$ such that $T,\Pi[\pi\mapsto t],i\models\varphi$,
  • $T,\Pi,i\models\forall\pi.\ \varphi$ if $T,\Pi[\pi\mapsto t],i\models\varphi$ for all $t\in T$,
  • $T,\Pi,i\models\exists\texttt{q}.\ \varphi$ if there exists a $t_{\texttt{q}}\in(2^{\{\texttt{q}\}})^{\omega}$ such that $T[\texttt{q}\mapsto t_{\texttt{q}}],\Pi,i\models\varphi$, and
  • $T,\Pi,i\models\forall\texttt{q}.\ \varphi$ if $T[\texttt{q}\mapsto t_{\texttt{q}}],\Pi,i\models\varphi$ for all $t_{\texttt{q}}\in(2^{\{\texttt{q}\}})^{\omega}$.
@@
While it is known that $\mathrm{Hyper\-QPTL}$ finite-state satisfiability is $\Sigma^{0}_{1}$-complete [11], i.e., complete for the recursively-enumerable languages, and $\mathrm{Hyper\-QPTL}$ model-checking is Tower-complete [11], the exact complexity of $\mathrm{Hyper\-QPTL}$ satisfiability is open: it is only known to be $\Sigma_{1}^{1}$-hard, as satisfiability is $\Sigma_{1}^{1}$-complete for its fragment $\mathrm{Hyper\-LTL}$ [7].
@@ 3.2 HyperQPTL Satisfiability
In this subsection, we study the $\mathrm{Hyper\-QPTL}$ satisfiability problem, which asks whether, for a given $\mathrm{Hyper\-QPTL}$ sentence $\varphi$, there is a nonempty $T$ such that $T\models\varphi$. We show that it is $\Sigma^{2}_{1}$-complete: intuitively, quantification over traces is equivalent to quantification over subsets of $\mathbb{N}$: a trace $t$ encodes the set $\{n\in\mathbb{N}\mid\texttt{p}\in t(n)\}$ and vice versa, where p is some designated proposition. Furthermore, existential quantification over sets of sets of natural numbers is equivalent to the selection of a set of traces at the meta-level of the satisfiability problem.
For this to be correct, we need to enforce that the model of a $\mathrm{Hyper\-QPTL}$ sentence contains enough traces to encode all subsets of $\mathbb{N}$. As usual, we use $\mathfrak{c}$ to denote the cardinality of the continuum, or equivalently, the cardinality of $(2^{\mathrm{AP}})^{\omega}$ for each finite $\mathrm{AP}$ and the cardinality of $2^{\mathbb{N}}$. As models of $\mathrm{Hyper\-QPTL}$ sentences are sets of traces, $\mathfrak{c}$ is a trivial upper bound on the size of models. It is straightforward to show that this upper bound is tight, which also implies that there are sentences whose models allow us to encode all subsets of $\mathbb{N}$.
@@
We need to show that every language in $\Sigma^{2}_{1}$ is polynomial-time reducible to $\mathrm{Hyper\-QPTL}$ satisfiability. To this end, we first present a polynomial-time computable function $f$ mapping sentences of the form $\varphi=\exists\mathcal{Y}_{1}\ldots\exists\mathcal{Y}_{k}.\ \psi$, with third-order variables $\mathcal{Y}_{j}$ and with $\psi$ only containing first-order and second-order quantifiers, to $\mathrm{Hyper\-QPTL}$ sentences $f(\varphi)$ such that $(\mathbb{N},+,\cdot,<,\in)\models\varphi$ if and only if $f(\varphi)$ is satisfiable. To this end, we fix $\mathrm{AP}=\{\texttt{x},\texttt{q},\texttt{m}_{1},\ldots,\texttt{m}_{k},\texttt{arg1},\texttt{arg2},\texttt{res},\texttt{add},\texttt{mult}\}$. We explain the roles of the different propositions along the way.
@@
$\theta_{\mathit{cons}}=\bigwedge\nolimits_{j=1}^{k}\forall\pi.\ \mathop{\mathbf{X}\vphantom{a}}\nolimits\mathop{\mathbf{G}\vphantom{a}}\nolimits\neg(\texttt{m}_{j})_{\pi}\wedge\forall\pi^{\prime}.\ \Big(\big(\mathop{\mathbf{G}\vphantom{a}}\nolimits(\texttt{x}_{\pi}\leftrightarrow\texttt{x}_{\pi^{\prime}})\big)\rightarrow\big((\texttt{m}_{j})_{\pi}\leftrightarrow(\texttt{m}_{j})_{\pi^{\prime}}\big)\Big)$
to express consistency: the first line expresses that only the first position of a trace may (but must not) be labeled with markers and the second line expresses that when two traces have the same truth value of x everywhere (i.e., they encode the same set), then they have the same marking.
So, every model $T$ of $\theta_{\mathit{\!all}}\wedge\theta_{\mathit{cons}}$ encodes every subset of $\mathbb{N}$ as a trace in $T$ and also $k$ sets of subsets of $\mathbb{N}$ via the (consistent) marking of traces, i.e., the existential quantification over a model in the satisfiability problem simulates the existential third-order quantification of the $\mathcal{Y}_{j}$.
Finally, Fortin et al. showed that addition and multiplication can be “implemented” in $\mathrm{Hyper\-LTL}$ [7]: Let $T_{(+,\cdot)}$ be the set of all traces $t\in(2^{\mathrm{AP}})^{\omega}$ such that there are unique $n_{1},n_{2},n_{3}\in\mathbb{N}$ with $\texttt{arg1}\in t(n_{1})$, $\texttt{arg2}\in t(n_{2})$, and $\texttt{res}\in t(n_{3})$, and either
@@
Now, we define for $\varphi=\exists\mathcal{Y}_{1}\ldots\exists\mathcal{Y}_{k}.\ \psi$
@@
  • For second-order variables $Y$, ${\mathit{hyp}}(\exists Y.\ \psi)=\exists\pi_{Y}.\ {\mathit{hyp}}(\psi)$ and ${\mathit{hyp}}(\forall Y.\ \psi)=\forall\pi_{Y}.\ {\mathit{hyp}}(\psi)$.
  • For first-order variables $y$, ${\mathit{hyp}}(\exists y.\ \psi)=\exists\pi_{y}.\ ((\neg\texttt{x}_{\pi_{y}})\operatorname{\mathbf{U}}(\texttt{x}_{\pi_{y}}\wedge\mathop{\mathbf{X}\vphantom{a}}\nolimits\mathop{\mathbf{G}\vphantom{a}}\nolimits\neg\texttt{x}_{\pi_{y}}))\land{\mathit{hyp}}(\psi)$ and ${\mathit{hyp}}(\forall y.\ \psi)=\forall\pi_{y}.\ ((\neg\texttt{x}_{\pi_{y}})\operatorname{\mathbf{U}}(\texttt{x}_{\pi_{y}}\wedge\mathop{\mathbf{X}\vphantom{a}}\nolimits\mathop{\mathbf{G}\vphantom{a}}\nolimits\neg\texttt{x}_{\pi_{y}}))\rightarrow{\mathit{hyp}}(\psi)$,